### v1

# Price of Uncertainty for Consensus Games

Abstract
Many game-theoretic models assume that players have access to accurate information, but uncertainty in observed data is frequently present in real-world settings. In this paper, we consider a model of uncertainty where adversarial perturbations of relative magnitude $1+\varepsilon$ are introduced to players’ observed costs. The effect of uncertainty on social cost is referred to as the price of uncertainty. We prove a tight bound on the price of uncertainty for consensus games of $\Theta(\varepsilon^{2}n^{2})$ for all $\varepsilon=\Omega\mathopen{}\left(n^{-1/4}\right)$. This improves a previous lower bound of $\Omega(\varepsilon^{3}n^{2})$ as well as a previous upper bound of $O(\varepsilon n^{2})$.
1 Introduction
If you are asked whether more of your friends like vanilla ice cream or chocolate ice cream, could you truly give the right answer with confidence? Even the best polling organizations sometimes fail to predict the outcomes of two-party majority elections. The truth is that uncertainty is everywhere and can impact our abilities to make rational decisions. However, most game-theoretic models assume that players have access to accurate information about interactions as well as their own costs. In the context of strategic interactions, it is interesting to see how uncertainty can affect agents’ behavior.
Balcan et al. (2009) consider how introducing small perturbations in observed costs in well-studied potential games can impact their behavior under improved-response and best-response dynamics. They coin the term price of uncertainty, and study it in a wide variety of potential games including set-covering games (Buchbinder et al., 2008), fair cost-sharing games (Anshelevich et al., 2008), job-scheduling games (Nisan et al., 2007), and consensus games. In some games, even extremely small perturbations can cause these dynamics to snowball out of control and move to states of much higher social cost, whereas in other games these dynamics are much more stable against large degrees of perturbation.
Consensus games are a special case of party affiliation games, which are in turn a class of potential games (Bhalgat et al., 2010). In a consensus game, each of $n$ players can choose between two parties: the white group and the red group. Players accrue a constant positive cost for interacting with those that chose the other side. Players are modeled as vertices of an undirected graph that can be colored either white or red, and a player’s cost is the number of neighbors with the other color, which we call bad neighbors. We will also call edges whose endpoints have different colors bad edges.
Starting from a ground state with social cost at least 1, the price of uncertainty is the maximum multiplicative factor by which the social cost can increase via uncertain best-response dynamics. The term uncertain here refers to the fact that a player can miscount their cost, which recall is the number of bad neighbors, by a multiplicative factor of at most $1+\varepsilon$ for a small uncertainty factor $\varepsilon>0$ that can depend on $n$. Note that we can assume $\varepsilon=\Omega\left(\frac{1}{n}\right)$, otherwise a player cannot miscount its number of bad neighbors and hence there is no uncertainty. Balcan et al. (2011) prove an upper bound of $O(\varepsilon n^{2})$ on the price of uncertainty for consensus games that holds for all $\varepsilon=\Omega\left(\frac{1}{n}\right)$ as well as a lower bound of $\Omega(\varepsilon^{3}n^{2})$ that holds in the domain $\varepsilon=\Omega\left(n^{-1/3}\right)$.
This paper obtains a lower bound for the price of uncertainty of consensus games of $\Omega\left(\varepsilon^{2}n^{2}\right)$ for all $\varepsilon=\tilde{\Omega}\left(n^{-1/2}\right)$. This is achieved using a construction that amplifies the snowball effect of uncertainty described by Balcan et al. (2011). Then, we will prove that the price of uncertainty for consensus games is also bounded by $O\left(\varepsilon^{2}n^{2}\right)$ for all $\varepsilon=\Omega\left(\varepsilon^{-1/4}\right)$ through the construction and analysis of a clever monovariant. These bounds coincide, so we conclude that the price of uncertainty for consensus games is $\Theta\left(\varepsilon^{2}n^{2}\right)$ for $\varepsilon=\Omega\left(n^{-1/4}\right)$, resolving an open problem proposed by Balcan et al. (2011).
2 Lower Bound
We first improve the lower bound for the price of uncertainty for consensus games. Furthermore, we do so in the expanded domain $\varepsilon=\tilde{\Omega}\left(n^{-1/2}\right)$. Here $\tilde{\Omega}$ hides factors that are logarithmic in $n$.
Theorem 1.
The price of uncertainty for consensus games is $\Omega(\varepsilon^{2}n^{2})$ for $\varepsilon=\tilde{\Omega}\left(n^{-1/2}\right)$.
To prove their lower bound, Balcan et al. (2009) used a carefully tuned gadget to create a snowball-like effect of uncertainty. To prove Theorem 1, we present a more complicated construction that amplifies the snowball effect even further.
2.1 Proof of Improved Lower Bound
Define an uncertain best response as a best response that strictly increases the social cost as a result of uncertainty. By definition, an uncertain best response must involve a color switch. One general idea is to construct a graph as a sequence of levels where each level contains slightly more vertices than the previous level. Starting from the first level, there are enough bad edges for all vertices in the first level to play uncertain best responses, which increases the number of bad edges by a multiplicative factor of $1+\varepsilon$. The fact that the vertices in the first level switched color implies that the vertices in the second level now have enough bad neighbors to play an uncertain best response, and this effect snowballs across the levels.
We now make the above idea rigorous. we define a boosting operation between layers of the following form. Let $\ell>1$ and suppose that level $\ell-1$ has $\frac{1}{1+\frac{\varepsilon}{2}}\cdot m$ vertices for some $m$, level $\ell$ has $m$ vertices, and level $\ell+1$ has $\left(1+\frac{\varepsilon}{2}\right)\cdot m$ vertices. Suppose that every vertex in level $\ell$ is connected to every vertex in level $\ell-1$ and level $\ell+1$, so the degree of every vertex in level $\ell$ is $d_{\ell}=\left(\frac{1}{1+\frac{\varepsilon}{2}}+(1+\frac{\varepsilon}{2})\right)\cdot m$. If the vertices in levels $\ell$ and $\ell+1$ are all white and the vertices in level $\ell-1$ are all red, then the edges between levels $\ell-1$ and $\ell$ are all bad. Since $(1+\varepsilon)\cdot\frac{1}{1+\frac{\varepsilon}{2}}\geq\frac{d_{\ell}}{2}$ for all $\varepsilon\leq 1$ then all vertices in level $\ell$ are allowed to play uncertain best responses in sequence, turning level $\ell$ all red. Note now that the number of bad edges between levels $\ell$ and $\ell+1$ is $\left(1+\frac{\varepsilon}{2}\right)\cdot m^{2}$, which is a $1+O(\varepsilon)$ multiplicative increase from the number of original bad edges $\frac{1}{1+\frac{\varepsilon}{2}}\cdot m^{2}$ from before this boosting operation.
We now present the construction. In order to implement the idea above, we must first find some way to create the first two layers of size $\frac{1}{\varepsilon}$ since a vertex must have degree at least $\frac{1}{\varepsilon}$ in order to play an uncertain best response. The initializer layers will have numbers of vertices
$\frac{1}{\varepsilon},1,\frac{1}{\varepsilon}+1,1,\ldots,\frac{2}{\varepsilon}-1,1,\frac{2}{\varepsilon},2,\frac{1}{\varepsilon},2,\frac{1}{\varepsilon}+1,2,\ldots,\frac{2}{\varepsilon}-1,2,\frac{2}{\varepsilon},4,\frac{1}{\varepsilon},\ldots,\frac{1}{\varepsilon},\frac{1}{\varepsilon}$
where the odd indexed layers cycle the consecutive integers from $\frac{1}{\varepsilon}$ to $\frac{2}{\varepsilon}$ and the numbers of vertices in even indexed layers increase by a factor of 2 whenever the odd indexed layer’s cycle repeats until the number reaches $\frac{1}{\varepsilon}$. Each vertex is connected to all vertices in the layer immediately after it except for vertices in odd indexed layers with $\frac{2}{\varepsilon}$ vertices, which are connected to the next layer to form a regular bipartite graph with the next layer where the vertices in the next layer are each connected to $\frac{1}{\varepsilon}$ vertices in the current layer. Note that the total number of initializer layers is $\frac{1}{\varepsilon}\cdot\log_{2}\frac{1}{\varepsilon}$ and each layer has at most $\frac{2}{\varepsilon}$ vertices, so the total number of vertices in the initializer layers is $O\left(\frac{1}{\varepsilon^{2}}\log\frac{1}{\varepsilon}\right)$. The initializer sequence has been constructed in such a way that that starting from a state in which the first layer is red and all other layers are white, the vertices in each layer can play uncertain best responses to change their color to red in sequence so that the last two layers form a $K_{\frac{1}{\varepsilon},\frac{1}{\varepsilon}}$ where the left side is red and the right half is white. Finally, we construct $\frac{1}{\varepsilon}$ white vertices in a parallel layer to the red left side of the $K_{\frac{1}{\varepsilon},\frac{1}{\varepsilon}}$ and connect each of them to every vertex in the right half of the $K_{\frac{1}{\varepsilon},\frac{1}{\varepsilon}}$, so that each vertex in the right half now has $\frac{1}{\varepsilon}$ red neighbors and $\frac{1}{\varepsilon}$ white neighbors on the left. Note that no new bad edges are added. We will now leave all layers before the current layer frozen, that is, no layer in a layer before the current layer will ever change color again.
After the current layer, construct a secondary set of layers with numbers of vertices
$\frac{1}{\varepsilon},1,\frac{1}{\varepsilon}+1,1,\ldots,\frac{2}{\varepsilon}-1,1,\frac{2}{\varepsilon},2,\frac{1}{\varepsilon},2,\frac{1}{\varepsilon}+1,2,\ldots,\frac{2}{\varepsilon}-1,2,\frac{2}{\varepsilon},4,\frac{1}{\varepsilon},\ldots,\frac{2}{\varepsilon},\frac{2}{\varepsilon}.$
Note that this sequence is almost identical to that of the initializer except we keep going until a $K_{\frac{2}{\varepsilon},\frac{2}{\varepsilon}}$ is formed by the last two layers. The first layer in the secondary set is the same as the last layer of the initializer, each vertex of which is connected with $\frac{1}{\varepsilon}$ frozen white vertices and $\frac{1}{\varepsilon}$ frozen red vertices. Note the following observation:
Proposition 2.
No matter what happens in future layers, each of the $\frac{1}{\varepsilon}$ vertices in the last layer of the initializer can freely change colors.
Proof.
Each vertex in this layer has total degree $\frac{2}{\varepsilon}+1$ and will always have at least $\frac{1}{\varepsilon}$ frozen neighbors of every color, which is enough to play an uncertain best response no matter what the current color is. ∎
It is our careful construction of the initializer to get to this point combined with Proposition 2 that is the key observation needed to improve the bound in Balcan et al. (2011). We now finish the construction. Note that the $K_{\frac{2}{\varepsilon},\frac{2}{\varepsilon}}$ produced at the end of the secondary set is a starting point from which we can perform the boosting operations as defined above. Here we add a layer of $\frac{2}{\varepsilon}+1$ vertices to form an initial $K_{\frac{2}{\varepsilon},\frac{2}{\varepsilon}+1}$ and then initialize $m$ to be $\frac{2}{\varepsilon}+1$. Starting from this layer, layers will increase in size by a constant multiplicative factor of $1+\frac{\varepsilon}{2}$ until we reach $c\cdot\varepsilon n$ vertices in the last layer for some small constant $c$ to be set later. The total number of vertices used in the boosting operations is $\Theta\left(\frac{c\cdot\varepsilon n}{1-\frac{1}{1+\frac{\varepsilon}{2}}}\right)=\Theta(cn)$. Combined with the number of vertices $O\left(\frac{1}{\varepsilon^{2}}\log\frac{1}{\varepsilon}\right)$ used in the initializer as well as the secondary set, we see that for $\varepsilon=\tilde{\Omega}\left(n^{-1/2}\right)$ and a constant $c$ independent of $\varepsilon$ and $n$ that the total number of vertices used is at most $n$, which means our construction has the right number of vertices.
How many bad edges does our construction yield? Starting from the first layer in the secondary set, we can turn each successive layer red in sequence so that the last two layers form a $K_{c\cdot\varepsilon n,c\cdot\varepsilon n}$ where the left side is red and the right half is white. However, this produces only $c^{2}\cdot\varepsilon^{2}n^{2}$ bad edges, which, recalling that we started with $\frac{1}{\varepsilon}$ bad edges, matches the bound $\Omega(\varepsilon^{3}n^{2})$ on the price of uncertainty in Balcan et al. (2011). It turns out we can produce many more bad edges in our construction by doing the following. Note that the last layer in our construction is white, and all layers between the frozen layers defined above and the last layer are now red. We now turn each vertex in the layer after the frozen layers white, which is a valid uncertain best response by Proposition 2. We can now turn each successive layer white in sequence until the last three layers are white, red, and white respectively. Repeatedly switching the colors of the vertices in the layer after the frozen layers and propagating uncertain best responses, we can reach a state in which the boosting operation layers alternate between white and red. Every edge between boosting operation layers is now bad, which makes $\Theta\left(\frac{c^{2}\cdot\varepsilon^{2}n^{2}}{1-\left(\frac{1}{1+\frac{\varepsilon}{2}}\right)^{2}}\right)=\Theta(\varepsilon n^{2})$ bad edges at the ending state. Recalling that we began with $\frac{1}{\varepsilon}$ bad edges, we have established that the price of uncertainty for consensus games is $\Omega(\varepsilon^{2}n^{2})$ in the domain $\varepsilon=\tilde{\Omega}\left(n^{-1/2}\right)$.
3 Upper Bound
We first provide a simple proof, omitted in Balcan et al. (2011), that the price of uncertainty for consensus games is $O(\varepsilon n^{2})$.
Theorem 3 ((Balcan et al., 2011, Theorem 2)).
The price of uncertainty for consensus games is $O(\varepsilon n^{2})$.
Proof.
We prove that in order to increase the number of bad edges, there must have been $\Omega\left(\frac{1}{\varepsilon}\right)$ bad edges at the start. Consider the first step at which the number of bad edges increases. Note that in order for a vertex’s uncertain best response to increase the number of bad edges, the vertex must think it has more bad neighbors than it actually has, hence the number of bad edges must be $\Omega\left(\frac{1}{\varepsilon}\right)$. Since this is the first step at which the number of bad edges increases, the number of bad edges at the start must be $\Omega\left(\frac{1}{\varepsilon}\right)$. Since there can be at most $\binom{n}{2}=O(n^{2})$ bad edges at any step, we conclude that the price of uncertainty is $O(\varepsilon n^{2})$. ∎
3.1 Ideas for improving the upper bound
The second main result of this paper is that the true price of uncertainty is $O(\varepsilon^{2}n^{2})$.
Theorem 4.
The price of uncertainty for consensus games is $O(\varepsilon^{2}n^{2})$ for $\varepsilon=\Omega\left(n^{-1/4}\right)$.
Generally, the goal of an adversary that wants to maximally blow up the number of bad edges can be thought of as trying to maximize the number of additional bad edges that each new vertex produces in the uncertain best-response process. Consider a time in which there are $m^{2}$ bad edges. The most “compact” way to store these bad edges, that is, uses the least number of vertices, is a bipartite graph $K_{m,m}$. If each of the $m$ vertices on one side of the bipartite graph plays an uncertain best-response, we can blow up the number of bad edges by a multiplicative factor of $(1+\varepsilon)$. However, this requires $(1+\varepsilon)m$ additional new vertices in order to “absorb” the new bad edges produced by the uncertain best-response. Hence to increase the number of bad edges from $m^{2}$ to $(1+\varepsilon)m^{2}$ requires roughly $(1+\varepsilon)m$ new vertices.
For example, following this heuristic for constant $\varepsilon$, that is $\varepsilon$ that is not a function of $n$, the last stage will have a $K_{\Theta(\varepsilon n),\Theta(\varepsilon n)}$. Let us be optimistic in the sense that we assume that the original bad edges at each stage, which become good after vertices play an uncertain bad response, can be turned bad again. Assuming most edges can eventually be turned bad again as is the case in Section 2, the total number of bad edges at the end is
$O\left(\varepsilon^{2}n^{2}+\frac{\varepsilon^{2}n^{2}}{1+\varepsilon}+\cdots\right)=O(\varepsilon n^{2}).$
Since there must be $\Omega\left(\frac{1}{\varepsilon}\right)$ bad edges at the start as in Theorem 3, we estimate the upper bound on price of uncertainty as $O(\varepsilon^{2}n^{2})$.
Motivated by the intuition in Section 3.1, it appears natural to track the number of bad edges at every step and try to establish an upper bound on its growth rate as new vertices play uncertain bad responses. The hope is that we can prove an upper bound on the growth rate in line with the rate described in Section 3.1. Unfortunately, this approach fails, as illustrated by the following example.
Example 5.
Consider a white vertex $V$ connected to $m$ white vertices and $m+1$ red vertices for $m=\Omega\left(\frac{1}{\varepsilon}\right)$. The $m$ white vertices are all connected to another white vertex $W$ and the $m+1$ red vertices are all connected to another red vertex $R$. We claim that we can have $V$ play $2m+1$ best responses in order to roughly double the number of bad edges while only operating on these $O(m)$ vertices:
  • Let $V$ play an uncertain best response that switches its color to red, resulting in $m+1$ bad edges.
  • Take the endpoint of one such bad edge adjacent to $W$ and switch its color so that there are now $m$ bad edges incident to $V$ and 1 bad edge incident to $W$.
  • Switch the color of $V$ again, resulting in $m+1$ bad edges. Take the endpoint of one such bad edge adjacent to $R$ and switch its color so that there are now $m$ bad edges incident to $V$, 1 bad edge incident to $W$, and 1 bad edge incident to $R$.
  • Repeat the above, at each step switching the color of a neighbor of $V$ whose color has not been switched before, until all neighbors of $V$ have had their colors switched once. There are now $m$ bad edges incident to $R$ and $m+1$ bad edges incident to $W$.
Note that Example 5 can be generalized to operate on a $K_{m,2m+1}$ where each vertex in the $m$-size part is white, $m$ of the vertices in the $(2m+1)$-size part is red, and $m+1$ of the vertices in the $(2m+1)$-size part is white to also roughly double the number of bad edges. The bad edge multiplicative growth rate of this process operating on $\Theta(m)$ vertices is 2, whereas in Section 3.1 it is $1+\varepsilon$. Hence using this analysis it is actually possible for there to be $\Omega(n^{2})$ bad edges at the last stage if operating on $m$ vertices can double the number of bad edges at each step. We thus cannot prove an upper bound of $O(\varepsilon n^{2})$ bad edges at the last stage using this strategy.
Instead, one clever idea is that we should focus not on the total number of bad edges at every step but rather consider the time step at which each vertex first plays an uncertain best response. This imposes a sequential order on the vertices from the first vertex to play an uncertain best response to the last. Recall that in order for a vertex with high degree to be able to play an uncertain best response, it must be the case that a large number of edges adjacent to this vertex must be bad. If we are to end up with $\Omega(\varepsilon n^{2})$ bad edges at the end, it must be the case that the later vertices in this sequence must have high degree on average. This requires a large number of existing bad edges to be generated from the vertices before the final vertices in the sequence, which in turn requires bad edges from the vertices before them, and so on.
Our proof of the $O(\varepsilon^{2}n^{2})$ upper bound analyzes the degrees of vertices in this sequence and shows that the sequence of degrees cannot increase “too fast” if vertices can only play $(1+\varepsilon)$-uncertain best responses, where we realize the notion of “too fast” as a clever monovariant.
3.2 Proof of Improved Upper Bound
We now give the formal proof of Theorem 4. It will be important to first establish a bound upon the starting number of bad edges.
Lemma 6.
In order to obtain a price of uncertainty on the order of $\Omega\left(\varepsilon^{2}n^{2}\right)$, the starting number of bad edges must be equal to $O\left(\varepsilon^{-2}\right)$.
Proof.
There can be at most $O\left(n^{2}\right)$ bad edges in a graph with $n$ vertices, so the starting number of bad edges must be
$\frac{O\left(n^{2}\right)}{\Omega\left(\varepsilon^{2}n^{2}\right)}=O\left(\varepsilon^{-2}\right).$
Let $G$ be the set of vertices of a graph upon which a consensus game is being played. Define $S_{0}$ as the set of vertices in $G$ that begin incident to at least one bad edge. Consider the set of vertices that begin without any bad edges and play at least one uncertain best response. They can be numbered in order of their first uncertain best response from first to last as $v_{1},v_{2},\dots,v_{m}$. Define a set $V$ as $\{v_{1},v_{2},\dots,v_{m}\}$. Finally, let $L$ be the set of vertices that begin without any bad edges and never play an uncertain best response. Notice that $S_{0}$, $V$, and $L$ are pairwise disjoint, and $S_{0}\cup V\cup L=G$.
For all integers $1\leq k\leq m$, let $S_{k}=S_{k-1}\cup\{v_{k}\}$.
Lemma 7.
Let $k$ be an integer between 0 and m1m-1. Before $v_{k+1}$ plays its first uncertain best response, every bad edge is incident to at least one vertex in $S_{k}$. Furthermore, every bad edge is always incident to at least one vertex in $S_{m}$.
Proof.
Before any uncertain best responses are played, all bad edges are incident to vertices in $S_{0}$ by definition. Similarly, at this point, edges between two vertices in $G\setminus S_{0}$ must all be good.
Let $0\leq j\leq m-1$ be an integer. Suppose that Lemma 7 holds when $k=j$. Then, any edge between two vertices in $G\setminus S_{j}$ is good. By construction, the first vertex outside of $S_{j}$ to play an uncertain best response is $v_{j+1}$. When this happens, the only edges to be affected will be those incident to $v_{j+1}$. However, since $v_{j+1}\in S_{j+1}$, every bad edge must now be incident to at least one vertex in $S_{j+1}$. Finally, as long as no vertices outside of $S_{j+1}$ change colors, every bad edge will remain incident to at least one vertex in $S_{j}$. By induction, Lemma 7 must hold for $k$ between 0 and m1m-1.
Finally, by construction, all vertices in $G\setminus S_{m}$ never play an uncertain best response. This means that any edges incident to 2 of these vertices will remain good, so Lemma 7 holds for $k=m$ as well. ∎
Lemma 8.
If $\varepsilon=\omega\left(n^{-1/2}\right)$, it is optimal to have $m=\Theta(n)$.
Proof.
Clearly, $m=|V|\leq|G|=n$, so $m=O(n)$, so the upper bound holds. We will now attack the lower bound.
By construction, $G$ is comprised of the two disjoint sets $S_{m}$ and $L$. By Lemma 7, any vertices in $L$ that do not share an edge with a vertex in $S_{m}$ can never be incident to a bad edge. This means that we can delete these vertices along with all of their incident edges, which will always increase the final price of uncertainty. By construction, prior to when $v_{1}$ plays its first uncertain best response, any edges incident to 2 vertices in $G\setminus S_{0}$ must be good. Since $L=G\setminus S_{m}\subset G\setminus S_{0}$ and since vertices in $L$ never change their colors, any edges incident to two vertices in $L$ must always be good. By removing all such edges, we will always increase the price of uncertainty.
We are now left with a new set of vertices $L^{\prime}\subset L$ where every edge incident to a vertex in $L^{\prime}$ is also incident to a vertex in $S_{m}\setminus S_{0}=V$. For $1\leq k\leq m$, $v_{k}$ can only play its first uncertain best response if its number of bad edges is at least its number of good edges multiplied by $(1+\varepsilon)^{-1}$. By Lemma 7, the number of bad edges incident to $v_{k}$ is at most $|S_{k-1}|\leq|S_{m}|$. Right before $v_{k}$ plays its first uncertain best response, all edges incident to $v_{k}$ and a vertex in $L^{\prime}$ must be good, so
$|S_{m}|^{2}\cdot(1+\varepsilon)^{2}$ $\geq|S_{m}|\cdot|L^{\prime}|$
$|S_{m}|\cdot(1+\varepsilon)^{2}$ $\geq|L^{\prime}|$
$\Theta\left(|S_{m}|\right)$ $=\Theta\left(L^{\prime}\right).$
Thus, the number of vertices in the graph has the same asymptotic growth as that of $S_{m}$. Recall that $|S_{m}|=|S_{0}|+m$. By Lemma 6, we must have $O\left(\varepsilon^{-2}\right)$ edges that start as bad, and the number of vertices that begin incident to a bad edge is at most twice this quantity. So, $|S_{0}|=O\left(\varepsilon^{-2}\right)$ as well.
$\Theta(n)$ $=|S_{m}|$
$\Theta(n)$ $=|S_{0}|+m$
$\Theta(n)$ $=O\left(\varepsilon^{-2}\right)+m$
Since $\varepsilon=\omega\left(n^{-1/2}\right)$, this becomes
$\Theta(n)$ $=o\left(n\right)+m$
$\Theta(n)$ $=m.$
With our construction, Lemma 7 allows us to contain the bad edges in the graph inside one of $S_{0},S_{1},\dots,S_{m}$. We will now present a combinatorial abstraction that will allow us to focus on the effect of the first uncertain best response played by vertices in $V$.
We can use $S_{k}$ to construct an infinite sequence $E_{k}$ of positive integers, finitely many of which are nonzero. Each nonzero element of $E_{k}$ will encode the number of edges incident to a vertex in $G\setminus S_{k}$ and any vertex in $S_{k}$, and each vertex in $G\setminus S_{k}$ will be encoded exactly once in such a manner. The rest of $E_{k}$ will be populated with zeros. Finally, to ensure that $E_{k}$ is unique, we will also include a requirement that $E_{k}$ must be sorted in descending order. For example, suppose that there existed 4 edges incident to a vertex in $S_{0}$ and some vertex $x_{1}$, and 3 edges incident to a vertex in $S_{0}$ and some vertex $x_{2}\neq x_{1}$. Also, suppose that there existed no other edges incident to a vertex in $S_{0}$ and a vertex in $G\setminus S_{0}$. Then, we would have
$E_{0}=\{4,3,0,0,0,\dots\}.$
Consider the state of the graph right before $v_{k}$ plays its first uncertain best response. By Lemma 7, all edges incident to $v_{k}$ and a vertex in $G\setminus S_{k-1}$ must be good. If we let $z$ be the number of edges incident to $v_{k}$ and a vertex in $S_{k-1}$, then the degree of $v_{k}$ cannot be greater than $z+\lfloor(1+\varepsilon)z\rfloor$. From this logic, going from $E_{k-1}$ to $E_{k}$ consists of deleting the element of $E_{k-1}$ corresponding to edges incident to a vertex in $S_{k-1}$ and $v_{k}$, which is $z$. We will now select $\lfloor(1+\varepsilon)z\rfloor$ elements of $E_{k-1}$ and increment them by $1$. This represents the new outgoing edges from $v_{k}$ to $G\setminus S_{k}$. Since any graph and sequence of uncertain best responses can be represented using the $E_{k}$ in such a manner, it suffices to study the behavior of these moves.
Definition 9.
A move on a sequence $E_{k}$ consists of the following steps:
  • Select some integer $z$ from $E_{k}$.
  • Delete one instance of $z$ from $E_{k}$.
  • Create $E_{k+1}$ by selecting $\lfloor z(1+\epsilon)\rfloor$ elements of $E$ and adding 1 to each of them.
For the rest of this proof, let $\sum E_{k}$ denote the sum of the elements of $E_{k}$.
Lemma 10.
If $\sum E_{m}=O\left(\varepsilon^{3}m^{2}|S_{0}|\right)$, then the final number of bad edges in the graph is $O\left(\varepsilon^{2}n^{2}|S_{0}|\right)$ for all $\varepsilon=\omega\left(n^{-1/2}\right)$.
Proof.
Let $T_{k}$ denote the number of edges that are incident to two vertices in $S_{k}$. Let $\Delta T_{k}$ denote $T_{k}-T_{k-1}$, and let $\Delta\sum E_{k}$ denote $(\sum E_{k})-(\sum E_{k-1})$. For every move that we play on an element $z$ in $E_{k}$, the sum $\sum E_{k}$ grows by $\lfloor z(1+\varepsilon)\rfloor-z$, and the number of edges that are contained inside of $S_{k}$ grows by $z$. This is because, by definition, $z$ denotes the number of edges that connect $S_{k}$ to the new vertex, and it is precisely these edges that are counted by $T_{k}$ and not $T_{k-1}$. We have
$\lfloor\varepsilon z\rfloor=\lfloor z(1+\varepsilon)\rfloor-z=\Delta\sum E_{k}$
$\varepsilon\cdot\Delta T_{k}-1=\varepsilon z-1<$ $\lfloor\varepsilon z\rfloor=\Delta\sum E_{k}$
$\varepsilon\cdot\Delta T_{k}-1<$ $\Delta\sum E_{k}$
By summing both sides from $k=1$ to $m$, we get
$\varepsilon\cdot\left(T_{m}-T_{0}\right)-m$ $\leq\sum E_{m}-\sum E_{0}.$
This lemma requires that $\sum E_{m}=O\left(\varepsilon^{3}m^{2}|S_{0}|\right)$.
$\varepsilon\cdot\left(T_{m}-T_{0}\right)-m$ $\leq O\left(\varepsilon^{3}m^{2}|S_{0}|\right)-\sum E_{0}$
By Lemma 8, this becomes
$\varepsilon\cdot\left(T_{m}-T_{0}\right)-\Theta(n)$ $\leq O\left(\varepsilon^{3}n^{2}|S_{0}|\right)-\sum E_{0}$
$\varepsilon T_{m}$ $\leq O\left(\varepsilon^{3}n^{2}|S_{0}|\right)-\sum E_{0}+\Theta(n)+\varepsilon T_{0}$
$\varepsilon T_{m}$ $\leq O\left(\varepsilon^{3}n^{2}|S_{0}|\right)+\Theta(n)+\varepsilon T_{0}$
$T_{m}$ $\leq O\left(\varepsilon^{2}n^{2}|S_{0}|\right)+\Theta\left(\varepsilon^{-1}n\right)+T_{0}.$
Since $\Omega\left(\varepsilon^{-1}\right)$ bad edges are required to play the first uncertain best response, we must have $S_{0}=\Omega\left(\varepsilon^{-1}\right)$. This means that $O\left(\varepsilon n^{2}\right)$ must also be $O\left(\varepsilon^{2}n^{2}|S_{0}|\right)$. If $\varepsilon=\omega\left(n^{-1/2}\right)$, then $O\left(\varepsilon n^{2}\right)$ will dominate $\Theta\left(\varepsilon^{-1}n\right)$.
$T_{m}$ $\leq O\left(\varepsilon^{2}n^{2}|S_{0}|\right)+T_{0}.$
Any vertex in $G$ is either in $S_{m}$, or $L$. This means that the only bad edges that are not counted by $T_{m}$ must be incident to a vertex in $S_{m}$ and a vertex in $L$. The edges that are incident to a vertex in $S_{m}$ and $G\setminus S_{m}=L$ are entirely described by $E_{m}$. Thus, the number of bad edges that exist between a vertex in $S_{m}$ and a vertex in $L$ is given by $\sum E_{m}$, which is $O\left(\varepsilon^{2}m^{2}\right)$. This is dominated by $T_{m}=O\left(\varepsilon n^{2}\right)+T_{0}$, so the asymptotic growth rate of the number of bad edges in the graph is given solely by that of $T_{m}$.
It remains to show that $T_{0}$ cannot grow faster than $O\left(\varepsilon^{2}n^{2}|S_{0}|\right)\geq\Theta(\varepsilon n^{2})$. By construction, every vertex in $S_{0}$ must begin incident to at least one bad edge, so the starting number of bad edges is at least $T_{0}/2=\Theta(T_{0})$. By Lemma 6, we start with $O\left(\varepsilon^{-2}\right)$ bad edges, so $T_{0}=O\left(\varepsilon^{-2}\right)$. This is also $O(\varepsilon n^{2})$ because $\varepsilon=\omega\left(n^{-1/2}\right)\geq\Theta\left(n^{-2/3}\right)$. ∎
Lemma 11.
The quantity $\sum E_{0}$ grows on the order of $O\left(m|S_{0}|\right)$.
Proof.
Since there are only $n$ vertices in the graph, every vertex in $S_{0}$ can have at most $n$ edges. Recall that $\sum E_{0}$ counts the number of edges incident to a vertex in $S_{0}$ and a vertex in $G\setminus S_{0}$. By Lemma 8, we have
$\sum E_{0}$ $=O\left(n|S_{0}|\right)$
$\sum E_{0}$ $=O\left(m|S_{0}|\right).$
We will now consider the time reversed process of playing moves on the $E_{k}$.
Definition 12.
A reversed move on a sequence $E_{k}$ consists of the following steps:
  • Select a list of positive elements $\{e_{1},e_{2},\dots,e_{\alpha}\}$ from $E_{k}$
  • Subtract one from each of the $e_{j}$ for $1\leq j\leq\alpha$
  • Calculate $z=\lceil\alpha/(1+\epsilon)\rceil$
  • Create $E_{k-1}$ by inserting $z$ into $E_{k}$ as a new element
Let $\sum E_{k}^{2}$ denote the sum of the squares of the elements of $E_{k}$. In order to investigate the asymptotic growth rate of $\sum E_{m}$, we wish to find the largest starting value of $\sum E_{m}$ for which it is possible to play $m$ reversed moves to get to a desired value of $\sum E_{0}$.
Our plan is to bound the behavior of $\sum E_{k}$ and $\sum E_{k}^{2}$ as reversed moves are played. Intuitively, playing a reversed move on many elements of $E_{k}$ would result in a large decrease in from $\sum E_{k}$ to $\sum E_{k-1}$. However, playing such a move would also result in the creation of a new, very large element in $E_{k-1}$. Since we can only ever subtract one from each element per move, having the sum $\sum E_{k-1}$ concentrated in a few large elements will make decreasing $\sum E$ more difficult. We consider $\sum E_{k}^{2}$ as a way to measure how ”spread out” the sequence $E_{k}$ is, and this particular quantity is preferable over other measures because of how the change from $\sum E_{k}^{2}$ to $\sum E_{k-1}^{2}$ can be cleanly bounded by an expression in terms of $\sum E_{k}$.
Consider some arbitrary sequence $E_{k}$. Let $\alpha$ denote the number of elements of $E$ that we play the current reversed game move on, and let these elements be denoted $e_{1},e_{2},\dots,e_{\alpha}$. By the rules of the reversed game, playing such a move will create a new element of size $z=\lceil\alpha/(1+\varepsilon)\rceil$. It is easy to see that the effect of such a move on $\sum E_{k}$ is
$\sum E_{k-1}=\sum E_{k}-\alpha+z.$
Next, we must have
$\sum E_{k-1}^{2}$ $=\sum E_{k}^{2}-\sum_{j=1}^{\alpha}\left(e_{j}^{2}-(e_{j}-1)^{2}\right)+z^{2}$
$=\sum E_{k}^{2}-2\sum_{j=1}^{\alpha}e_{j}+\alpha+z^{2}.$
Since $e_{1},e_{2},\dots,e_{\alpha}$ are all elements of $E_{k}$, their sum must be bounded by $\sum E_{k}$. By plugging this into our expression for the change in $\sum E^{2}$, we get
$\sum E_{k-1}^{2}-\sum E_{k}^{2}$
$={}$ $-2\sum_{j=1}^{\alpha}e_{j}+\alpha+z^{2}$
$\geq{}$ $-2\sum_{j=1}^{\alpha}e_{j}+z^{2}$
$\geq{}$ $-2\sum E_{k}+\left\lceil\frac{\alpha}{1+\varepsilon}\right\rceil^{2}$
$\geq{}$ $-2\sum E_{k}+(1-\varepsilon)^{2}\alpha^{2}$ (1)
Lemma 13.
It is always optimal to take the lower bound in (1) for the change in $\sum E^{2}$.
Proof.
Notice that neither our change in $\sum E$ nor our bound (1) on the change in $\sum E^{2}$ depend on $\sum E_{k}^{2}$. If we make $\sum E_{k}^{2}$ as small as possible for a fixed $\sum E_{k}$, the future options we will have for $\alpha$ will not change. Furthermore, we only care that $\sum E_{0}^{2}$ ends up being smaller than $\left(\sum E_{0}\right)^{2}$. This means that allowing $\sum E_{k-1}^{2}$ to attain its lower bound will never result in more moves being required. There are no further issues because there is no need that intermediate values can be represented by actual sequences; there is no required minimum value for $\sum E_{k}^{2}$ for any integer $0\leq k\leq m-1$. ∎
We can also obtain a simpler lower bound for the change in $\sum E$.
$\sum E_{k-1}-\sum E_{k}$
$={}$ $-\alpha+z$
$={}$ $\left\lceil\frac{\alpha}{1+\varepsilon}\right\rceil-\alpha$
$\geq{}$ $(1-\varepsilon)\alpha-\alpha$
$={}$ $-\varepsilon\alpha.$ (2)
Lemma 14.
It is always optimal to take the lower bound in (2) for the change in $\sum E$.
Proof.
If we allow $\alpha\in\mathbb{R}$, still with $\alpha\geq 1$, it is always possible to choose a new $\alpha^{\prime}$ such that
$\varepsilon\alpha^{\prime}=\left\lceil\frac{\alpha}{1+\varepsilon}\right\rceil-\alpha.$
The RHS of the previous equation is clearly nonpositive, so $\alpha^{\prime}$ must be nonpositive as well. We can also bound the size of $\alpha^{\prime}$ with
$-\varepsilon\alpha^{\prime}$ $=\left\lceil\frac{\alpha}{1+\varepsilon}\right\rceil-\alpha$
$1-\varepsilon\alpha^{\prime}$ $\geq 1+\frac{\alpha}{1+\varepsilon}-\alpha$
$-\varepsilon\alpha^{\prime}$ $\geq\frac{-\varepsilon\alpha}{1+\varepsilon}$
$(1+\varepsilon)\alpha^{\prime}$ $\leq\alpha$
$\alpha^{\prime}$ $<\alpha.$
By (1), if we take the lower bound for the change in $\sum E_{k}^{2}$ using $\alpha^{\prime}$, the value of $\sum E_{k-1}^{2}$ is equal to or smaller than if we had used $\alpha$ instead. By the same logic that was used to prove Lemma 13, it is always helpful to have a lower $\sum E^{2}$. ∎
With Lemma 13 and Lemma 14, we now have a strictly stronger set of actions on $\sum E$ and $\sum E^{2}$ that we can perform with each move.
Definition 15.
A strong reversed move that uses $\alpha$ on the pair of values $\sum E_{k}$ and $\sum E_{k}^{2}$ is given by
$\sum E_{k-1}$ $=\sum E_{k}-\varepsilon\alpha$
$\sum E_{k-1}^{2}$ $=\sum E_{k}^{2}-2\sum E_{k}+(1-\varepsilon)^{2}\alpha^{2}.$
For integers $1\leq k\leq m$, let $\alpha_{k}$ denote the value of $\alpha$ used in the strong reversed move to get from $\left(\sum E_{k},\sum E_{k}^{2}\right)$ to $\left(\sum E_{k-1},\sum E_{k-1}^{2}\right)$. Then,
$\sum E_{m}$ $=\sum E_{0}+\varepsilon\alpha_{1}+\varepsilon\alpha_{2}+\cdots+\varepsilon\alpha_{m}$
$\sum E_{m}$ $=\sum E_{0}+\varepsilon\sum_{j=1}^{m}\alpha_{j}.$ (3)
The same process is a little more complicated for $\sum E_{m}^{2}$.
$\sum E_{m}^{2}$ $=\sum E_{0}^{2}+2\sum_{j=1}^{m}\left(\sum E_{j}\right)-(1-\varepsilon)^{2}\sum_{j=1}^{m}\alpha_{j}^{2}$
We can rewrite the $\sum E_{j}$ by plugging in (3).
$\sum E_{m}^{2}$ $=\sum E_{0}^{2}+2\sum_{j=1}^{m}\left(\sum E_{j}\right)-(1-\varepsilon)^{2}\sum_{j=1}^{m}\alpha_{j}^{2}$
$\sum E_{m}^{2}$ $=\sum E_{0}^{2}+2\sum_{j=1}^{m}\left(\sum E_{0}+\varepsilon\sum_{i=1}^{j}\alpha_{i}\right)-(1-\varepsilon)^{2}\sum_{j=1}^{m}\alpha_{j}^{2}$
$\sum E_{m}^{2}$ $=\sum E_{0}^{2}+2m\sum E_{0}+2\varepsilon\sum_{j=1}^{m}(m-j+1)\alpha_{j}-(1-\varepsilon)^{2}\sum_{j=1}^{m}\alpha_{j}^{2}$
$0$ $=\sum E_{0}^{2}+2m\sum E_{0}-\sum E_{m}^{2}+2\varepsilon\sum_{j=1}^{m}(m-j+1)\alpha_{j}-(1-\varepsilon)^{2}\sum_{j=1}^{m}\alpha_{j}^{2}$
We have $\sum E_{m}^{2}\geq\sum E_{m}$. Furthermore, due to (3), $\sum E_{m}\geq\varepsilon\sum_{j=1}^{m}\alpha_{j}$. We can use these two inequalities to replace $\sum E_{m}^{2}$.
$0$ $\leq\sum E_{0}^{2}+2m\sum E_{0}-\varepsilon\sum_{j=1}^{m}\alpha_{j}+2\varepsilon\sum_{j=1}^{m}(m-j+1)\alpha_{j}-(1-\varepsilon)^{2}\sum_{j=1}^{m}\alpha_{j}^{2}$
$0$ $\leq\sum E_{0}^{2}+2m\sum E_{0}-\varepsilon\sum_{j=1}^{m}\alpha_{j}+2\varepsilon\sum_{j=1}^{m}(m)\alpha_{j}-(1-\varepsilon)^{2}\sum_{j=1}^{m}\alpha_{j}^{2}$
$0$ $\leq\sum E_{0}^{2}+2m\sum E_{0}+\varepsilon(2m-1)\sum_{j=1}^{m}\alpha_{j}-(1-\varepsilon)^{2}\sum_{j=1}^{m}\alpha_{j}^{2}$ (4)
By the Cauchy-Schwarz Inequality, we have
$\left(\sum_{j=1}^{m}1\right)\left(\sum_{j=1}^{m}\alpha_{j}^{2}\right)$ $\geq\left(\sum_{j=1}^{m}\alpha_{j}\right)^{2}$
$\sum_{j=1}^{m}\alpha_{j}^{2}$ $\geq m^{-1}\left(\sum_{j=1}^{m}\alpha_{j}\right)^{2}.$
We can then plug this into (4):
$0$ $\leq\sum E_{0}^{2}+2m\sum E_{0}+\varepsilon(2m-1)\sum_{j=1}^{m}\alpha_{j}-(1-\varepsilon)^{2}m^{-1}\left(\sum_{j=1}^{m}\alpha_{j}\right)^{2}$
$0$ $\geq(1-\varepsilon)^{2}m^{-1}\left(\sum_{j=1}^{m}\alpha_{j}\right)^{2}-\varepsilon(2m-1)\sum_{j=1}^{m}\alpha_{j}-\left(\sum E_{0}^{2}+2m\sum E_{0}\right).$
We are now left with a quadratic in the sum $\sum_{j=1}^{m}\alpha_{j}$. Since the leading coefficient is positive and the quadratic is on the lesser side of the inequality, $\sum_{j=1}^{m}\alpha_{j}$ must be less than the larger root of this quadratic. This larger root is obtained through a tedious application of the quadratic formula:
$\sum_{j=1}^{m}\alpha_{j}$ $\leq\frac{\varepsilon(2m-1)+\sqrt{\varepsilon^{2}(2m-1)^{2}+4(1-\varepsilon)^{2}m^{-1}\left(\sum E_{0}^{2}+2m\sum E_{0}\right)}}{(1-\varepsilon)^{2}m^{-1}}.$ (5)
Lemma 16.
If $\varepsilon=\Omega\left(n^{-1/4}\right)$, then $\sum_{j=1}^{m}\alpha_{j}$ is bounded by $O\left(\varepsilon^{2}m^{2}|S_{0}|\right)$.
Proof.
If we substitute asymptotic behavior into the RHS of (5), we get
$\sum_{j=1}^{m}\alpha_{j}$ $\leq\frac{\Theta(\varepsilon m)+\sqrt{\Theta\left(\varepsilon^{2}m^{2}\right)+\Theta\left(m^{-1}\sum E_{0}^{2}\right)+\Theta\left(\sum E_{0}\right)}}{\Theta\left(m^{-1}\right)}.$
Notice that we must have $\sum E_{0}^{2}\leq\left(\sum E_{0}\right)$. By Lemma 11, we can rewrite the previous inequality as follows:
$\sum_{j=1}^{m}\alpha_{j}$ $\leq\frac{\Theta(\varepsilon m)+\sqrt{\Theta\left(\varepsilon^{2}m^{2}\right)+\Theta\left(m|S_{0}|^{2}\right)+\Theta\left(m|S_{0}|\right)}}{\Theta\left(m^{-1}\right)}$
It is clear that the $\Theta\left(m|S_{0}|^{2}\right)$ term dominates the $\Theta\left(m|S_{0}|\right)$ term. We now have the following:
$\sum_{j=1}^{m}\alpha_{j}$ $\leq\frac{\Theta(\varepsilon m)+\sqrt{\Theta\left(\varepsilon^{2}m^{2}\right)+\Theta\left(m|S_{0}|^{2}\right)}}{\Theta\left(m^{-1}\right)}.$ (6)
Depending on which one of the two terms inside of the square root dominate, there are two cases that we have to tackle.
Case 16.1.
$\Theta\left(\varepsilon^{2}m^{2}\right)$ dominates $\Theta\left(m|S_{0}|^{2}\right)$.
In this case, we can rewrite (6) as follows:
$\sum_{j=1}^{m}\alpha_{j}$ $\leq\frac{\Theta(\varepsilon m)+\sqrt{\Theta\left(\varepsilon^{2}m^{2}\right)}}{\Theta\left(m^{-1}\right)}$
$\sum_{j=1}^{m}\alpha_{j}$ $\leq\frac{\Theta(\varepsilon m)}{\Theta\left(m^{-1}\right)}$
$\sum_{j=1}^{m}\alpha_{j}$ $=O\left(\varepsilon m^{2}\right).$
Recall that $|S_{0}|=\Omega\left(\varepsilon^{-1}\right)$. This means that $O\left(\varepsilon^{2}m^{2}|S_{0}|\right)$ must include $O\left(\varepsilon m^{2}\right)$, and we are done with this case.
Case 16.2.
$\Theta\left(m|S_{0}|^{2}\right)$ dominates $\Theta\left(\varepsilon^{2}m^{2}\right)$.
In this case, we can rewrite (6) as follows:
$\sum_{j=1}^{m}\alpha_{j}$ $\leq\frac{\sqrt{\Theta(\varepsilon^{2}m^{2})}+\sqrt{\Theta\left(\varepsilon^{2}m^{2}\right)+\Theta\left(m|S_{0}|^{2}\right)}}{\Theta\left(m^{-1}\right)}$
$\sum_{j=1}^{m}\alpha_{j}$ $\leq\frac{\sqrt{\Theta(\varepsilon^{2}m^{2})}+\sqrt{\Theta\left(m|S_{0}|^{2}\right)}}{\Theta\left(m^{-1}\right)}$
$\sum_{j=1}^{m}\alpha_{j}$ $\leq\frac{\sqrt{\Theta\left(m|S_{0}|^{2}\right)}}{\Theta\left(m^{-1}\right)}$
$\sum_{j=1}^{m}\alpha_{j}$ $=O\left(m^{3/2}|S_{0}|\right).$
By Lemma 8, $\varepsilon=\Omega\left(n^{-1/4}\right)$ is equivalent to $\varepsilon=\Omega\left(m^{-1/4}\right)$.
$\sum_{j=1}^{m}\alpha_{j}$ $=O\left(\varepsilon^{2}\cdot\varepsilon^{-2}m^{3/2}|S_{0}|\right)$
$\sum_{j=1}^{m}\alpha_{j}$ $=O\left(\varepsilon^{2}\left(m^{-1/4}\right)^{-2}m^{3/2}|S_{0}|\right)$
$\sum_{j=1}^{m}\alpha_{j}$ $=O\left(\varepsilon^{2}m^{2}|S_{0}|\right)$
The result of Lemma 16 can be plugged into (3) with Lemma 11 to obtain a bound on $\sum E_{m}$. Since we are only interested in $\varepsilon=\Omega\left(n^{-1/4}\right)\geq\Theta\left(n^{-1/3}\right)$, we have
$\sum E_{m}$ $=\sum E_{0}+\varepsilon\sum_{j=1}^{m}\alpha_{j}$
$\sum E_{m}$ $=O\left(m|S_{0}|\right)+O\left(\varepsilon^{3}m^{2}|S_{0}|\right)$
$\sum E_{m}$ $=O\left(\varepsilon^{3}m^{2}|S_{0}|\right).$ (7)
By (7) and Lemma 10, the final number of bad edges in the graph must be $O\left(\varepsilon^{2}n^{2}|S_{0}|\right)$. By construction, every vertex in $S_{0}$ begins with at least one bad edge, so we must begin with $\Omega\left(|S_{0}|\right)$ bad edges. We obtain the following upper bound on the price of uncertainty for consensus games when $\varepsilon=\Omega\left(n^{-1/4}\right)$:
$\frac{O\left(\varepsilon^{2}n^{2}|S_{0}|\right)}{\Omega\left(|S_{0}|\right)}=O\left(\varepsilon^{2}n^{2}\right).$
This finishes the proof of Theorem 4.